OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3)

OpenStudy (lgbasallote):

distribute first \[\implies -2x + 6 \ge 5 - x - 3\] do you know how to solve for x now?

OpenStudy (anonymous):

noo

OpenStudy (lgbasallote):

do you know how to solve for x in \[-2x + 6 = 5 - x - 3\]??

OpenStudy (anonymous):

-3x + 6=5-3 -3x+ 9=5 -3x=-4?

OpenStudy (anonymous):

hello??

OpenStudy (lgbasallote):

you add x

OpenStudy (lgbasallote):

-2x + x = -x not -3x

OpenStudy (lgbasallote):

so that should be -x = -4 got it?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

2(x + 5) > 8x – 8 help with this bro